1. Suppose we want to determine the (binomial) probability (p) of getting 6 heads in 13 flips of a 2-sided coin. Using Table 2 in Appendix B of the text, what values of n, x, and p would we use to look up this probability, and what would be the probability?
This is what I think I know:
n = 11
x = 6
p = .50q = 1-p = .50
I think this is the forumula I use:
11/(11-6)6 (.50)^2(.50)^5 =
I can't figure out any more or even if what I have done is correct. I need some line for line, number to number detail so I can really learn and understand this please.

Answer by stanbon(75887)

Suppose we want to determine the (binomial) probability (p) of getting 6 heads in 13 flips of a 2-sided coin.
---
The 6 heads can show up is 13C6 = 1716 mutually exclusive ways.
---
Each of those ways has probability (1/2)^6*(1/2)^7 = 1/2^13 = 1/8192
====
Because the 1716 ways to succeed are mutually exclusive you should
add all 1716 of those separate probabilities to get the final probability.
------
You could add but it's easier to just multiply 1716*(1/8192) = 0.2095
---
In general, if there are n trials, and the probability of success
on each trial is "p", the probability of "k" successes is as follows:
P(x = k) = nCk*p^k*q^(n-k)
==============================
Cheers,
stan H.
===============